# Ion Temperature Anisotropy Limits from Magnetic Curvature Scattering in Magnetotail Reconnection Jets

### v1

@@ Acknowledgements.
We thank the MMS team for data access and support. L.R. acknowledges support from the Royal Swedish Academy of Sciences grant AST2024-0015, the Knut and Alice Wallenberg Foundation (Dnr. 2022.0087), and the Swedish National Space Agency grant 192/20. CN acknowledges support from the Swedish National Space Agency (SNSA) under Grant 2022-00121, and the Research Council of Norway under contract 300865. The PIC-HC simulation was performed on resources provided by Sigma2 – the National Infrastructure for High-Performance Computing and Data Storage in Norway—X.A. is supported by NASA grant No. 80NSSC22K1634. X.A. acknowledges high-performance computing support from Derecho (https://doi.org/10.5065/qx9a-pg09), provided by NCAR’s Computational and Information Systems Laboratory and sponsored by the National Science Foundation [derecho]. A.V.A. and S.R.K. are supported by NASA grants No. 80NSSC23K0658, 80NSSC25K0074, and 80NSSC22K0752.

### v2

@@ Acknowledgements.
We thank the MMS team for data access and support. L.R. acknowledges support from the Royal Swedish Academy of Sciences grant AST2024-0015, the Knut and Alice Wallenberg Foundation (Dnr. 2022.0087), and the Swedish National Space Agency grant 192/20. CN acknowledges support from the Swedish National Space Agency (SNSA) under Grant 2022-00121, and the Research Council of Norway under contract 300865. The PIC-HC simulation was performed on resources provided by Sigma2 – the National Infrastructure for High-Performance Computing and Data Storage in Norway. X.A. is supported by NASA grant No. 80NSSC22K1634. X.A. acknowledges high-performance computing support from Derecho (https://doi.org/10.5065/qx9a-pg09), provided by NCAR’s Computational and Information Systems Laboratory and sponsored by the National Science Foundation [24]. A.V.A. and S.R.K. are supported by NASA grants No. 80NSSC23K0658, 80NSSC25K0074, and 80NSSC22K0752.